### v2

@@ 6.1. Method
Participants. To ensure a diverse area of expertise, we recruited 14 participants via snowball sampling. Participants were screened using the same demographic questionnaire from the formative study. We followed best practices from the HCI Guidelines for Gender Equity and Inclusivity to collect gender-related information. Participants ($\mu=22.6$ years old, $\sigma=2.6$ years), denoted as E01-E14, represented 11 different areas of expertise and varying levels of familiarity in robotics and demonstration collection (Table A). The participants’ familiarity with robotics ranged from unfamiliar to familiar: three people were unfamiliar, six people were somewhat unfamiliar, three people had neutral familiarity, and two people were familiar. Regarding their familiarity with the demonstration collection, eight people were unfamiliar, three people were somewhat unfamiliar, two people had neutral familiarity, and one person was somewhat familiar. Each participant performed one to three tasks, resulting in a total of 30 collections (10 collections per task). None of the participants participated in the formative study.
@@
Procedure. The in-person laboratory study took place on a university campus, and each task lasted 40 minutes. The consent form was taken in person. Because we were interested in candid reactions, we did not tell participants that we created EVE. Each participant was compensated with food for the laboratory study.
We began by providing a short tutorial for all the collection interfaces (i.e., kinesthetic teaching, teleoperation with an HTC Vive VR controller, AR2-D2, and EVE) and informed them of the goal of the task(s). We did not demonstrate how to accomplish the task(s) because we did not want to influence the trajectory collected by the participants. After a 30-second practice session before each collection, participants were asked to complete the task in three minutes. We recorded each collection’s task success rate and remaining time (three minutes - collection time). We allowed the participant to retry the collection if a collision occurred and the collection time had not reached three minutes. After the participants had completed the collection with all the collection interfaces, we asked them to complete the SUS survey and a form to rank the motion intent communication, user enjoyment, and overall preference for all collection interfaces.
@@
Analysis. We analyzed the results from the SUS survey and ranking forms with the same method used for the formative study result analysis. See Table 2 for a summary of quantitative results from the post-task questionnaires. We calculated the overall task success rate and the mean and standard deviation of the remaining time for each collection interface.
@@ 6.2. Findings
EVE achieved the highest task success rate and performed comparably to kinesthetic teaching in task completion time, system usability, motion intent communication, user enjoyment, and overall user preference ($mean_{p}=0.30$). Kinesthetic teaching attained the highest SUS score and ranked the highest in motion intent, user enjoyment, and overall preference. However, there was no statistically significant difference between EVE and kinesthetic teaching, suggesting that EVE may be a viable substitute for collecting trajectory data. This might be because of the ability to clearly inspect the real world before moving the robot, a feature not present in AR2-D2 and teleoperation. Our results confirm the findings of previous work (Wrede et al., 2013; Mandlekar et al., 2018), which showed that kinesthetic teaching outperforms VR controllers in all measurements. EVE outperformed AR2-D2 in both demonstration collection success rate and remaining time. In all tasks performed by the participants, EVE surpassed AR2-D2 in SUS score, motion intent, enjoyment, and overall preference ranking.
Figure 5. The mean success rate (%) for the evaluation user study, which encompassed three tasks with 10 trials each ($D=10$), indicates that the EVE achieved a 100% success rate across all tasks. This performance surpasses that of the second-highest interface, kinesthetic teaching, by an average margin of 10% across the tasks.
Figure 6. The average remaining time in seconds and its standard deviation taken to successfully collected one demonstration for each tasks. EVE’s performance in terms of time efficiency is comparable to that of kinesthetic teaching, with a narrow average difference of 5.1 seconds across three tasks.
@@
Table 2. Mean and standard deviation of usability, motion intent communication, enjoyment, and overall preference for different demonstration collection interfaces. Usability is rated on a scale from 0 to 100, with higher scores indicating better usability. Motion intent communication, enjoyment, and overall preference are ranked from 1 to 4, with lower ranks indicating better performance. One asterisk (*) indicates statistical significance ($p<0.05$) when compared to the EVE system.
Below are some notable comments from users based on their observations of different collection interfaces used in three tasks during the evaluation study:
Toggle Switch. This task involves operating an augmented robot gripper to navigate around potential obstacles and accurately activate a switch. In kinesthetic teaching, users appreciated the method’s simplicity and precise effector control, facilitating movement to the desired location. However, this system does not support the modification of previously established waypoints. For teleoperation and AR2-D2, many users reported failures, attributing them to collisions with scene distractors and the systems’ inability to provide a comprehensive scene view due to robotic arm obstructions. Specifically, teleoperation users struggled with accurately positioning the robot end-effector to engage the switch, hindered by a lack of haptic feedback. Conversely, in our EVE system, users found the AR-KT features significantly improved control during the final positioning of the gripper, and the invisible robot mode provided a clear view to assess potential collisions with distractors.
Sort Food. This task required users to control a robot to perform a series of long-horizon actions, including picking up and placing multiple objects. In kinesthetic teaching, users found it straightforward to maneuver the robot for pick-and-place actions, praising its intuitiveness and ease of use. However, in teleoperation, the robot’s obstruction of the end-effector’s view complicated precise control, leading to difficulty in achieving the ideal grasp pose and frequent minor errors, often resulting in objects being in ungraspable states. In AR2-D2, users struggled due to inadequate visualization of potential collisions between the robot arm and table distractors. Conversely, the EVE system was well-received, with users appreciating the path history for tracking long-horizon actions and the redo feature for correcting waypoints. The AR-KT functionality was highlighted for its precise control, allowing for accurate object placement and enhanced gripper manipulation.
Sweep Table. In this task, users needed precise control to grasp onto the sponge and execute a series of sweeping actions. Kinesthetic teaching was favored for its simplicity, allowing users to manipulate the robot’s end-effector directly. For the teleoperation system, users encountered difficulties due to latency and the precision required to control the end-effector with a VR controller. They reported struggling to maintain a consistent direction of movement when controlling the end-effector, often leading to collisions between the end-effector and the table. With AR2-D2, the challenge was also in using fixed forward control to linearly move the end-effector along one axis, compounded by unreliable tracking of the desired gripper state. Conversely, the EVE system provided enhanced precision through its AR-KT features, offering accurate gripper control that ensured the gripper maintained the state desired by the user.
@@
In this section, we reflect on our findings from the development and evaluation of EVE. We also discuss future opportunities for research exploring AR-based demonstration collection interfaces for robotics.
@@ 7.1. Implications
Personalized Robot Training in Everyday Environments. By allowing users to easily capture videos and manipulate objects in various locations outside of traditional laboratory settings, EVE enables individuals to expose robots to a wide range of objects and environments that are relevant to their specific needs and preferences. As the cost of robot hardware decreases and household robots become more prevalent, EVE puts the power of robot training directly into the hands of users, allowing them to collect personalized demonstrations for specific household tasks and fine-tune their robots to perform these tasks effectively in their unique home environments.
@@
Integrating with Other Types of Robots. EVE allows users to quickly adapt the system to work with different robot platforms. Because each component of the program is separated, by replacing the robot prefab in Unity with a different robot model, such as the UR5e, users can collect demonstrations with a new AR robot. Furthermore, since the waypoint data collected by EVE is a standard input to multiple motion planners, EVE enables the training of manipulation policies for different robots without modifying the collected data. This flexibility empowers individuals to extend the applicability of EVE to a wide range of robotic systems, making robot training accessible to a broader audience.
7.2. Limitations
Multi-View Data Collection. EVE is limited to single-viewpoint data collection to maintain accurate hand-eye calibration between the robot’s actions and the corresponding 3D voxel representation, which is essential for successful real-world deployment. However, multi-view data collection could be helpful in scenarios where objects are occluded from one perspective but visible from another, or when the demonstrator needs to maintain a more comfortable posture by collecting trajectories from different angles. Future work could explore the use of simultaneous localization and mapping (SLAM) techniques (Durrant-Whyte and Bailey, 2006) to track the iPad LiDAR’s pose and build a 3D map of the workspace, allowing the demonstrator to collect trajectories relative to this 3D map.
Bimanual Robotic Manipulation. Many real-world tasks, such as washing dishes, ironing clothes, and assembling furniture, require bimanual manipulation. However, collecting demonstrations for bimanual tasks using EVE is challenging due to the inability to press buttons during collection. Future research could explore the use of mixed reality head-mounted displays for demonstration collection. With the increasing availability of novel head-mounted displays, collections using these devices might become more affordable and ubiquitous. Alternative input modalities, such as gaze and speech, could be used to trigger different functions instead of clicking buttons.
@@ 7.2. Limitations
Collection for Mobile Robot. The real world also involves tasks that require mobile robots, such as cleaning the floor and watering plants. Extending EVE to support demonstration collection for mobile robots presents an exciting direction for future work. To achieve this, researchers can integrate a 3D model representing the mobile robot into Unity and develop functionalities to control the robot’s movement through intuitive touch gestures. The system can then leverage the device’s camera and inertial measurement unit (IMU) sensors to construct a detailed map of the environment while simultaneously tracking the robot’s pose, enabling data collection for mobile robots.
@@
Hardware. The current system’s hand-tracking capabilities are not fully optimized due to hardware constraints. The LiDAR technology employed in the iPad occasionally results in inaccuracies in the end effector’s movement and insufficient anchoring of the robot base on the surface. To address this issue and improve tracking accuracy and robustness, future work can implement sensor fusion techniques by combining data from multiple sensors, such as the iPad’s camera, inertial measurement unit (IMU), and LiDAR scanner.
@@
We introduced EVE, a system that democratizes robot data collection with human-centric AR visualizations. Informed by our formative study ($N=10$), EVE enables end users to collect, change, and verify valuable trajectory data to train robot manipulation for personalized tasks. Our evaluation user study ($N=14,D=30$) demonstrated improved task success rate, completion time, system usability, motion intent communication, user enjoyment, and user preference against state-of-the-art collection interfaces. We hope EVE will catalyze future work in AR-based demonstration collection for robotics and expedite the incorporation of robots into daily lives.
@@ 10 F10 - Computer Science Neutral Somewhat unfamiliar Assemble car components
Table 3. Participant demographics in formative and evaluation study. In total, there were 24 participants overall (#). f denotes the participant IDs for the formative study, e denotes the participant IDs for the evaluation study.

### v3

@@ 6.1. Method
Participants. To ensure a diverse area of expertise, we recruited $14$ participants via snowball sampling. Participants were screened using the same demographic questionnaire from the formative study. To reduce bias, we ensured that none of the participants participated in the formative study. We followed best practices from the HCI Guidelines for Gender Equity and Inclusivity (Scheuerman et al., 2020) to collect gender-related information. Participants ($\mu=22.6$ years old, $\sigma=2.6$ years), denoted as E01-E14, represented 11 different areas of expertise and varying levels of familiarity in robotics and demonstration collection (Table A). The participants’ familiarity with robotics ranged from unfamiliar to familiar: three people were unfamiliar, six people were somewhat unfamiliar, three people had neutral familiarity, and two people were familiar. Regarding their familiarity with the demonstration collection, eight people were unfamiliar, three people were somewhat unfamiliar, two people had neutral familiarity, and one person was somewhat familiar. $11$ participants identified as male, while three participants identified as female.
@@
Procedure. The laboratory study was conducted on a university campus, with each session lasting $40$ minutes. Consent forms were collected in person. To ensure candid reactions, participants were not informed that we developed EVE. Participants were compensated with food for their participation in the study.
Each session began with a brief tutorial on all the collection interfaces, including kinesthetic teaching, teleoperation with an HTC Vive VR controller, AR2-D2, and EVE. Participants were informed about the goals of the tasks but were not shown how to accomplish them to avoid bias. After a $30$-second practice session for each interface, participants began the task collection.
A collection was deemed successful if the participant completed the task without any collisions within three minutes, excluding setup times. If a collision occurred and the three-minute limit had not been reached, participants were allowed to retry. For kinesthetic teaching and teleoperation, a collision was defined as the robot hitting objects that were not the target objects. For AR2-D2 and EVE, collisions were manually assessed by the study proctor by checking if any part of the AR robot intersected with physical objects during collection and task replay. This assessment strategy was chosen as the alternative would require expensive hardware setup with multiple cameras and constructing a depth map is computationally expensive for iPad.
Optimal data collection for AR2-D2 and EVE required that the distance between the AR robot and the scene match that of the physical robot and its surroundings. This is a common assumption in robotics research due to the fixed camera location relative to the robot. To ensure data accuracy, we marked a point on the table and positioned the AR robot at the same spot before each collection. future work could utilize physical reference markers to calibrate the camera and the collected data, thereby removing this assumption.
Success rates and remaining time for each task were recorded for all interfaces. The remaining time was used instead of total time to maintain consistency with the representation of success rates, where higher values indicate better performance. Data collection time for each interface was limited to three minutes, as it is challenging to collect a successful demonstration using a virtual reality controller for teleoperation.
Upon completing the task collection with all interfaces, participants filled out the SUS survey and a form ranking motion intent communication, user enjoyment, and overall preference for the interfaces. We aimed to measure $10$ collection attempts for each task. To prevent fatigue from using four interfaces per task, participants could stop at any time, resulting in each participant collecting between one and three tasks. All participants completed a NASA TLX survey post-study to assess the workload. Finally, we explained to the participant how the participant’s data would be used to investigate the design for AR-based demonstration collection interfaces and answered any relevant questions.
@@
Analysis. We analyzed the results from the SUS survey and ranking forms with the same method used for the formative study result analysis. See Table 3 for a summary of the quantitative results from the post-task questionnaires. Additionally, we calculated the overall task success rate and determined the mean and standard deviation for the remaining time and the NASA TLX survey.
@@ 6.2. Findings
EVE achieved the highest task success rate and performed comparably to kinesthetic teaching in terms of task completion time, system usability, motion intent communication, user enjoyment, and overall user preference ($mean_{p}=0.30$). As expected, Kinesthetic teaching attained the highest SUS score and was ranked highest in motion intent, user enjoyment, and overall preference. However, there was no statistically significant difference between EVE and kinesthetic teaching, suggesting that EVE may be a viable alternative for collecting trajectory data. This could be due to the ability to clearly inspect the environment during collection and the intuitive control of guiding the robot with the hand (AR-KT). Our results support previous findings (Wrede et al., 2013; Akgün and Subramanian, 2011), which showed that kinesthetic teaching outperforms VR controllers in accuracy, speed, and user preference. EVE outperformed AR2-D2 in all measurements. The NASA TLX survey post-study indicated a low to medium workload for all users after the collection ($\mu=2.57$, $\sigma=2.45$), further reinforcing the validity of our findings.
Figure 4. The mean task success rate (%) for all interfaces in the evaluation user study, which included three tasks with 10 trials each, indicates that EVE achieved the highest success rate across all tasks.
[…]
Bar graphs titled “Task Success Percentage” showing the three tabletop tasks in the evaluation study on the X axis and the mean success rate (%) of the four collection interfaces on the Y axis.
Figure 5. The mean and the standard deviation of the remaining time (seconds) for successfully completing one demonstration for each task. EVE performed comparably to kinesthetic teaching, with an average difference of 5.1 seconds across the three tasks.
[…]
Bar graphs titled “Task Remaining Time” showing the three tabletop tasks in the evaluation study on the X axis and the mean and standard deviation of the remaining time (sec) of the four collection interfaces on the Y axis. The standard deviation of the Teleoperation via VR controller for the sort food task is missing because only one person succeeded in the task collection.
@@
Table 3. EVE performed comparably to kinesthetic teaching in usability, motion intent communication, user enjoyment, and user preference. The mean and standard deviation of all measurements are shown. Usability is rated on a scale from 0 to 100, with higher scores indicating better usability. Motion intent communication, user enjoyment, and user preference are ranked from 1 to 4, with lower ranks indicating better performance. One asterisk (*) indicates statistical significance ($p<0.05$) when compared to the EVE system.
[…]
Mean and standard deviation of usability, motion intent, enjoyment, and preference of Kinesthetic Teaching, Teleoperation via VR controller, AR2-D2, and EVE.
Below are observations and feedback on using different collection interfaces for the tasks:
Toggle switch. This task involves navigating the AR robot’s end effector around physical obstacles and accurately pressing the switch. With kinesthetic teaching, users appreciated the method’s simplicity and the precise control of the end effector. However, some users needed to rotate the robot’s body during data collection to avoid collisions, which increased manual effort. Teleoperation with a VR controller proved challenging, with most users failing to collect successful demonstrations due to the unintuitive controls and the difficulty in accurately perceiving the switch’s position from the point cloud representation of the scene displayed on a 2D monitor. With AR2-D2, users frequently failed due to difficulties in interpreting the 3D position of the AR robot, resulting in demonstrations where the end effector touched points near the switch rather than the switch itself. Conversely, all users successfully collected demonstrations with EVE. Most users utilized the AR-KT feature, aided by hand position projection, to guide the AR robot to the switch.
Sort food. This is a long-horizon task requiring users to control a robot to pick up and place multiple objects. With kinesthetic teaching, many users appreciated the precise control of the gripper for pick-and-place actions. In contrast, teleoperation posed challenges, as most users struggled to pick up food due to positional discrepancies between the point cloud representation and the actual scene. With AR2-D2, users found it difficult to control the gripper for grasping and releasing objects. Additionally, many users failed to collect successful demonstrations because they assumed the AR robot would automatically avoid physical obstacles to reach the specified waypoint. Conversely, all users successfully collected demonstrations with EVE. Most users rotated the iPad camera to prevent collisions between the AR robot’s body and the blue cup.
Sweep table. This task required users to control a robot to grab a sponge and perform a series of sweeping actions. Many users preferred kinesthetic teaching for its simple control, allowing them to directly manipulate the robot’s end effector and receive haptic feedback when the end effector contacted the table surface. With the teleoperation system, users faced challenges due to latency and the precision needed to control the end effector with a VR controller. They struggled to maintain a consistent direction of movement, often resulting in collisions between the end effector and the table. With AR2-D2, users found it difficult to move the robot’s end effector horizontally, as specifying waypoints with the same vertical distance from the table surface was challenging. With EVE, most users favored using AR-KT to specify waypoints, as they found it easier to maintain the hand’s height during a continuous movement.
6.3. Real world deployment
To assess the effectiveness of the collected demonstrations, we trained robot policies for the toggle switch task using data gathered by AR2-D2 and EVE.
Training procedure. Like AR2-D2, EVE collects RGB-D data and waypoint coordinates during the data collection process. After collecting the data, we used Perceiver-Actor (PerAct) (Shridhar et al., 2022) to train a transformer-based language-guided behavior cloning policy. PerAct takes a 3D voxel observation and a language goal (v, l) as inputs and produces discretized outputs for the translation, rotation, and gripper state of the end-effector. These outputs, in conjunction with a motion planner, enable the execution of tasks specified by the language goal. We trained the policy using $6$ demonstrations over $30,000$ iterations for each interface.
Results. The trained policies were evaluated with $30$ task rollouts for the toggle switch task. The policy trained with EVE-collected data achieved a $30\%$ success rate, compared to a $20\%$ success rate with AR2-D2-collected data. We emphasize that the toggle switch task is more challenging than the button press task from the AR2-D2 paper, as it requires millimeter-level precision for the robot’s end-effector to touch a specific point on the switch to turn on a light bulb. The higher success rate with EVE-collected data may be attributed to the improved depth perception during collection, allowing users to adjust waypoints for more accurate trajectories. We have included a video demonstrating a successful rollout of the toggle switch task and the 3D voxel observations generated with data collected using AR2-D2 and EVE on our website [^2] https://junwang0510.github.io/EVE/.
@@ 7. Discussion
In this section, we reflect on insights from the development and evaluation of EVE. We also discuss future opportunities for research exploring AR-based demonstration collection interfaces for robotics.
@@ 7.1. Implications
EVE enables personalized robot training in everyday environments. EVE empowers users to gather demonstrations in diverse settings beyond traditional laboratories, exposing robots to a variety of objects and environments tailored to their specific needs. As robot hardware becomes more affordable and household robots more common, EVE puts the power of robot training directly into the hands of users, allowing them to collect personalized demonstrations for specific household tasks.
@@
EVE is compatible with various robot platforms and motion planners. EVE enables users to quickly adapt the system to different robot platforms. Since each program component is modular, users can gather data for another type of robot by replacing the robot prefab in Unity with another model, such as the UR5 robot. Additionally, since EVE collects waypoint data in a standard format compatible with common motion planners, it facilitates the training of manipulation policies for different robots without needing to modify the collected data. This adaptability extends EVE’s applicability to a wide range of robotic systems, making robot training more accessible to a broader audience.
7.2. Limitations and future work
Bimanual robotic manipulation. Many real-world tasks, such as washing dishes, ironing clothes, and assembling furniture, require the use of both hands. However, EVE currently does not support bimanual manipulations, as users need to press buttons during the collection process and it is uncomfortable to collect demonstrations with an iPad positioned in front of the body. Future research could address this limitation by exploring the use of AR head-mounted displays. This approach would allow users to control both robot arms with their hands while using voice commands to trigger specific operations. Additionally, these technologies enable the implementation of better depth cues, such as realistic shadows and dynamic lighting, to improve the user’s ability to interact with the AR environment accurately. However, implementing bimanual controls may present additional challenges, such as accurately tracking both hands during occlusion.
@@ 7.2. Limitations and future work
Collection for mobile navigation robots. The real world also involves tasks that require mobile robots, such as cleaning the floor and watering plants. Extending EVE to support demonstration collection for mobile robots is a promising direction for future research. This can be achieved by integrating a 3D model of the mobile robot into Unity and developing intuitive gesture controls for its movement. For instance, users could draw the desired trajectory for the robot on an iPad, similar to the TrajectoryFly feature in PinpointFly (Chen et al., 2021). The system could also utilize the device’s camera and inertial measurement unit (IMU) sensors to create a detailed map of the environment while tracking the robot’s pose, thereby enabling effective data collection for mobile robots.
Multi-view data collection. EVE is limited to single-viewpoint data collection to maintain accurate hand-eye calibration between the robot’s actions and the corresponding 3D voxel representation, which is essential for successful real-world deployment. However, multi-view data collection could be helpful in scenarios where objects are occluded from one perspective but visible from another, or when the demonstrator needs to maintain a more comfortable posture by collecting trajectories from different angles. Future work could explore the use of SLAM techniques (Durrant-Whyte and Bailey, 2006) to track the iPad LiDAR’s pose and build a 3D map of the workspace, allowing the demonstrator to collect trajectories relative to this 3D map.
Dynamic tasks. To enhance EVE’s capability to perform tasks in unstructured environments, we integrated motion planning algorithms such as Fast-RRT and RRT* by leveraging LiDAR-generated octree maps. However, the system failed to operate in real-time due to the iPad’s limited computational power. Future research could experiment with various edge devices and develop more efficient motion planners.
@@
Hardware. We acknowledge that the reliance on specific hardware (iOS devices with AR capabilities) might limit the user base. Future work could evaluate EVE on different devices (e.g., Android devices, 3D glasses) or in less ideal task environments to test the generalizability of our design and findings. Additionally, the iPad’s LiDAR camera occasionally results in inaccuracies in the end effector’s movement and insufficient anchoring of the robot base on the surface. To address this issue and improve tracking accuracy and robustness, future work could implement sensor fusion techniques by combining data from multiple sensors, such as the iPad’s camera, IMU, and other cameras.
@@ 8. Conclusion
We present EVE, an iOS application that democratizes robot data collection through human-centric AR visualizations. EVE empowers users to collect, correct, and inspect trajectory data to train robots for personalized tasks. In our evaluation user study, EVE outperformed three state-of-the-art interfaces in terms of success rate and matched the performance of kinesthetic teaching in completion time, usability, motion intent communication, enjoyment, and user preference. We hope EVE will catalyze future work in AR-based demonstration collection for robotics and expedite the incorporation of robots into daily lives.
Acknowledgements.
We are grateful to Abhimanyu Saighal for his assistance in preprocessing the collected demonstrations for training robot policies. We also thank Sebastin Santy and Zixian Ma for their valuable feedback on the paper.
@@ 10 F10 - Computer Science Neutral Somewhat unfamiliar Assemble car components
Table 4. Participant demographics in formative and evaluation studies. There were a total of 24 participants overall. f denotes participant IDs for the formative study, and e denotes participant IDs for the evaluation study.
[…]
Participant demographics in formative and evaluation study. There were 24 participants overall (#), 10 for the formative study (f denotes the participant IDs) and 14 for the evaluation study. We recorded the area of expertise, familiarity with robotics (5-point Likert scale), familiarity with demonstration collection (5-point Likert scale), and preferred tasks for the robot.